OpenStudy (vortish):

x/7 + 3x/14 less or equal to 4/7

OpenStudy (vortish):

just trying to figure out how the steps go do not want the answer just the steps to get there

OpenStudy (bahrom7893):

multiply everything by 14: 2x+3x<=8 5x<=8 divide everything by 5: x<=8/5

OpenStudy (vortish):

what i dont understand is where you get the 2x from

OpenStudy (bahrom7893):

(x/7) * 14 = 14x/7 = 2*7x/7=2x

OpenStudy (vortish):

ok that makes sense now tanks